Assume an economy has 2 firms. Firm A produces wheat and firm B produces bread. In a given year, firm A produces 50,000 bushels of wheat, sells 20,000 bushels of wheat to firm B at $3 per bushel, export 25,000 bushels of wheat at $3 per bushel and stores the rest as inventory. Firm A pays $50,000 in wages to consumers. Firm B produces 50,000 loaves of bread and sells all of it to domestic consumer at $2 per loaf. Firms B pays consumers $20,000 in wages. The economy also imports and consumes 15,000 loaves of bread for a $1 per loaf. (a) Calculate Net Exports. (b) Find Investment. (c) Calculate GDP.
Answer #1

a) Net Exports = Exports-Imports

Exports = 25000 × 3 = $75000

Imports = 15000 × 1 = $15000

Net Exports = $75000 - $15000 = $60000

b) Investment = Inventory = 5000 bushels of wheat = 5000 × 3 = $15000.

c) calculating GDP by value added method.

Firm A : 50000×3 =$150000

Firm B : (50000×2) - (20000×3) = $40000. $60000 is subtracted because the wheat sold by firm A is an input to firm B.

Total GDP = $150000 + $40000 =$190000.

If any query regarding the last part, please comment. Thank you.